Current System Status

Please see below for current and past incidents.

If you have an issue which you believe isn’t affected by a current incident, then please contact our Support Team on 03451 283458.

SYSTEM STATUS
Fri 24 September 2021

We experienced a hardware issue on our core that has subsequently been resolved. The majority of client ccts re-routed correctly as per design with a small number of clients however, requiring additional manual intervention.
This issue is now resolved however please contact Support should you still be experiencing problems.

Fri, 24 Sep 2021 09:15

We are now seeing the network and associated client devices recover and will continue to monitor to ensure stability. An RFO will be posted in due course

Fri, 24 Sep 2021 09:00

Investigations are ongoing

Fri, 24 Sep 2021 07:30

We are aware of an issue that is currently affecting a small number of clients. This is presently under investigation and further updates will be posted as they become available, We apologise for the disruption in the interim

 

Thu, 17 Jun 2021 19:31:42

We can now confirm resolution to this incident and all services are operational

Services affected Thursday 17 June – 16:30

We are seeing the majority of services as recovered now.
For those clients still experiencing issues we would advise an equipment reboot if not already done. Services should still be considered to be at risk however until we receive full resolution details

We are currently aware of major network incident affecting BT Broadband services. This is presently under investigation and further updates will be posted here when received

Services affected Tues 23 March – 15:00-16:00

We were aware of various reachability issues on the internet, that manifested themselves as intermittent service unavailability.
The root cause of this was an issue at the London Internet Exchange which is a traffic exchange point for most of the UK ISPs. The issue withdrew ISPs networks, and then reintroduced them, which had the effect of rapidly changing paths to destination traffic.
Securus has resilient connections to the whole internet and subsequently withdrew our peering to the London Internet Exchange to stabilise traffic.
From this point we saw stable traffic on our network. We reintroduced the London Internet Exchange at about 9pm Tues 23 March and traffic has remained stable. We continue to monitor this through our various systems.

13:36 14 Oct 2020
Engineers have completed their investigation and believe the issue to now be resolved. We apologise for any inconvenience caused.

13:05 14 Oct 2020
We are starting to see services recover. As yet, we are still awaiting confirmation of details relating to the cause and as such, the current status will remain intact until such time as we can confirm resolution

12:40 14 Oct 2020admin.php?page=custom-dashboard.phpadmin.php?page=custom-dashboard.php
We are aware of a number of clients affected by a recent service outage and are presently liaising for more detail which will be posted here in due course.

PAST INCIDENTS . . .

A small number of clients may be experiencing connectivity issues as a result of a Major service outage affecting some of our suppliers. The issue is under investigation and further updates to follow once available

Thu, 27 Aug 2020 18:57:52
We now believe all issues to have been resoladmin.php?page=custom-dashboard.phpved. the disruption was an unprecedented fire issue in a datacentre

Thu, 27 Aug 2020
inbound voice services are currently down for Securus and for some customers. For any issues, please email support@securuscomms.co.uk

August 18, 2020 at 22:57
We are advised that all services are now restored. All customers are now fed from the new infrastructure and at full redundancy. An RFO will be distributed in due course when full details are available. The incident will now be resolved as all systems are available with no further issues reported. Monitoring will remain in place to ensure stability.

August 18, 2020 at 19:30
We are now seeing the majority of services as recovered. We will continue to monitor the few remaining locations affected and advise once operational

August 18, 2020 at 16:15
services have been further restored to more cadmin.php?page=custom-dashboard.phpustomers and increasing numbers of those affected are now operational. Engineers continue to work towards restoring services in full

August 18, 2020 at 15:15
An increasing number of services are reported and seen to recover. Engineers continue to work towards restoring services in full

August 18, 2020 at 14:12:22
services have been further restored to more customers and increasing numbers of those affected are now operational. Engineers continue to work to all customers by migrating to the newly installed and commissioned infrastructure

August 18, 2020 at 13:00
Engineering reports that services have been further restored for several more customers and the team continues to work towards restoring services to all customers by migrating to the newly installed and commissioned infrastructure

August 18, 2020 at 11:55
Update: services have been reinstated for several customers and engineering teams continue to work towards restoring services to all customers by migrating to the newly installed and commissioned infrastructure. Further updates to follow.

August 18, 2020 at 11:00
Engineers have advised that their team have begun restoring power to affected devices. Unfortunately, at present there remains no estimate resolution time.
We will continue to provide updates as soon as they come through.

10:50
Services remain impacted
Investigation and work to resolve is ongoing

Client services are presently impacted due to an ongoing issue at a central Data Centre. Latest advisory confirms engineers are working to resolve. Further updates to follow . . .

11:20 – 26 Feb 2020
Please be advised we are currently experiencing issues with an upstream provider. A workaround is in place to resolve however clients should be aware of potential further disruption to services. Investigations are underway and further update will be posted in due course.. We apologise for the disruption to services at this time.

11:58 – 26 Feb 2020
We are seeing recovery across the Estate. Clients still experiencing connection issues are recommended to reboot their internet router

13:10 – 26 Feb 2020
Investigation into root cause is still underway.

15:15 – 26 Feb 2020
The issue has been identified as upstream supplier hardware. This has now been resolved and services have been restored.

None – Apr 20, 09:00 BST

READY TO WORK WITH US? THESE PEOPLE DO...